# Revisiting institutional punishment in the $N$-person prisoner’s dilemma

### v1

@@ I Introduction
The usual evolutionary game theoretical framework for studying cooperation is the $N$-person prisoner’s dilemma Fox_1978 . In a play group of $N>1$ individuals, each individual can decide whether to contribute a fixed amount $c>0$ to the common pool (public goods). This amount is multiplied by a factor of $r>1$ and then divided among the $N-1$ other players. Following the terminology of the $2$-person prisoner’s dilemma, we refer to individuals that contribute to the common pool as cooperators and those that do not as defectors. Clearly, in this scenario cooperators are doomed to extinction unless there is some positive assortment between them Hamilton_1975 , which is best illustrated by the green-beard effect Hamilton_1964 , or repeated interactions that allow the emergence of reputation and reciprocity Wang_2023 ; Quan_2024 ; Han_2024 ; Lai_2024 .
@@
The rest of the paper is organized as follows. In Section II, we present a variant of the $N$-person prisoner’s dilemma that more realistically models both the maintenance and the operation of sanctioning institutions. In particular, in this section we present the instantaneous payoffs of the three strategies that players can adopt, namely cooperators, defectors, and institutional punishers, and describe the stochastic imitation dynamics that govern the evolution of strategy frequencies for finite populations. In Section III we study the game for infinitely large populations using the replicator equations framework: we find the equilibrium solutions and the conditions for local stability, and explore the dynamics by numerically solving the replicator equations and representing the orbits in the simplex. In Section IV we study the finite population version of the game, using Monte Carlo simulations to implement the stochastic imitation dynamics and finite-size scaling techniques to infer fixation probabilities and mean time to fixation in the limit of very large populations. Finally, in Section V we summarize our results and present some concluding remarks.
@@ II The model
Consider a well-mixed population of finite size $M$, composed of $X$ cooperators, $Y$ defectors, and $Z$ institutional punishers, such that $X+Y+Z=M$. Since we are only considering institutional punishers here, we will refer to these individuals simply as punishers. The population is well mixed in the sense that every individual can interact with every other individual in the population. At each time step $\delta t$, a focal individual $i$ is randomly selected. The play group of $i$ is formed by randomly selecting $N-1\geq 1$ other individuals without replacement from the remaining $M-1$ individuals in the population.
@@
Once the focal individual’s payoff $f_{i}$ is determined, a model individual $j\neq i$ is randomly selected along with her play group, which may include the focal individual $i$ or other members of $i$’s play group. The payoff $f_{j}$ of the model individual is obtained using the same procedure described above. If $f_{j}>f_{i}$, then the focal individual $i$ switches to the strategy of the model individual $j$ with probability
@@
where $\Delta f_{\max}$ is the maximum possible payoff difference that guarantees that this ratio is less than or equal to 1. If $f_{j}\leq f_{i}$, then the focal individual $i$ keeps her strategy. In our variant of the NPD, we have
@@
which is the difference between the payoff obtained by a single defector in a group of cooperators (i.e., $rc$) and the payoff obtained by a single punisher in a group of defectors (i.e., $-c-\gamma$). This completes the time step $\delta t$ of the stochastic dynamics, and so the time is updated to $t=t+\delta t$. As usual in such an asynchronous update scheme, we choose the time increment to be $\delta t=1/M$, so that during the increment from $t$ to $t+1$ exactly $M$, though not necessarily different, individuals are selected as focal individuals. The particular choice of switching probability given in Eq. (4) produces the replicator equation in the limit $M\to\infty$ Fontanari_2024b (see Traulsen_2005 ; Sandholm_2010 for other switching probabilities that produce the replicator equation in the large population limit).
@@ III The replicator equation formulation
where $\bar{\pi}=x\pi^{C}+y\pi^{D}+z\pi^{P}$ is the average payoff of the population and $\Delta f_{\max}$ is given by Eq. (5). Although we could eliminate the factor $1/\Delta f_{\max}$ by a proper rescaling of $t$ and thus obtain the standard form of the replicator equation, keeping it facilitates comparison with the finite population simulations presented in Section IV. We refer the reader to Ref. Hofbauer_1998 for a thorough introduction to the replicator equation and to Ref. Fontanari_2024b for the derivation of the replicator equation as the deterministic limit of the stochastic imitation dynamics introduced in Section II. Since $x+y+z=1$ and the expected payoffs of the different strategies depend only on $x$ and $z$, henceforth we will eliminate $y$ in favor of these variables. The resulting equations become

### v2

@@ I Introduction
The usual evolutionary game theoretical framework for studying cooperation is the $N$-person prisoner’s dilemma Fox_1978 . In a play group of $N>1$ individuals, each individual can decide whether to contribute a fixed amount $c>0$ to the common pool (public goods). This amount is multiplied by a factor of $r>1$ and then divided among the $N-1$ other players. Following the terminology of the $2$-person prisoner’s dilemma, we refer to individuals that contribute to the common pool as cooperators and those that do not as defectors. Clearly, in this scenario cooperators are doomed to extinction unless there is some positive assortment between them Hamilton_1975 , which is best illustrated by the green-beard effect Hamilton_1964 , or repeated interactions that allow the emergence of reputation and reciprocity Wang_2023 ; Lai_2024 .
@@
The remaining sections are organized as follows. In Section II, we present a variant of the $N$-person prisoner’s dilemma that more realistically models both the maintenance and the operation of sanctioning institutions. In particular, in this section we present the instantaneous payoffs of the three strategies that players can adopt, namely cooperators, defectors, and institutional punishers, and describe the stochastic imitation dynamics that govern the evolution of strategy frequencies for finite populations. In Section III we study the game for infinitely large populations using the replicator equations framework: we find the equilibrium solutions and the conditions for local stability, and explore the dynamics by numerically solving the replicator equations and representing the orbits in the simplex. In Section IV we study the finite population version of the game, using Monte Carlo simulations to implement the stochastic imitation dynamics and finite-size scaling techniques to infer fixation probabilities and mean time to fixation in the limit of very large populations. In Section V, we recapitulate our main findings and present some closing observations.
@@ II The model
Consider a well-mixed population of finite size $M$, composed of $X$ cooperators, $Y$ defectors, and $Z$ institutional punishers, such that $X+Y+Z=M$. Since we are only considering institutional punishers here, we will refer to these individuals simply as punishers. A focal individual $i$ is randomly chosen at each time step $\delta t$. To determine her payoff $f_{i}$, we form her play group by randomly choosing another $N-1\geq 1$ individuals from the remaining $M-1$ individuals in the population without replacement.
@@
Once the payoff of the focal individual $f_{i}$ has been determined, we need to select a model individual $j\neq i$ in order to implement the imitation process. The model individual and her play group are selected randomly and the payoff $f_{j}$ is determined as done for the focal individual. The key component of the imitation dynamics is that individuals will only imitate their more successful peers, which means that the focal individual will not change her strategy in the case that $f_{j}\leq f_{i}$. However, when $f_{j}>f_{i}$, the probability that focal individual $i$ switches to the strategy of the model individual $j$ is
@@
where $\Delta f_{\max}$ is chosen so as to guarantee that this probability is not greater than 1. In our variant of the NPD, we have
@@
which is the difference between the payoff obtained by a single defector in a group of cooperators (i.e., $rc$) and the payoff obtained by a single punisher in a group of defectors (i.e., $-c-\gamma$). The update (or not) of the focal individual concludes the time step $\delta t$ and the time variable $t$ is increased accordingly. The particular choice of switching probability given in Eq. (4) produces the replicator equation in the $M\to\infty$ limit Fontanari_2024b (see Traulsen_2005 ; Sandholm_2010 for other switching probabilities that produce the replicator equation in the large-population limit), provided that the time step is set to $\delta t=1/M$.
@@ III The replicator equation formulation
where $\bar{\pi}=x\pi^{C}+y\pi^{D}+z\pi^{P}$ is the average payoff of the population and $\Delta f_{\max}$ is given by Eq. (5). Although we could eliminate the factor $1/\Delta f_{\max}$ by a proper rescaling of $t$ and thus obtain the standard form of the replicator equation, keeping it facilitates comparison with the finite population simulations presented in Section IV. As already mentioned, these equations are derived as the infinite population limit of the imitation dynamics described in Section II (see Fontanari_2024b ). Since $x+y+z=1$ and the expected payoffs of the different strategies depend only on $x$ and $z$, henceforth we will eliminate $y$ in favor of these variables. The resulting equations become